Crane Realty Corporation purchased a tract of unimproved land for $52,000. This land was improved and subdivided into building lots at an additional cost of $30,000. These building lots were all of the same size but owing to differences in location were offered for sale at different prices as follows Group No. of Lots Price per Lot $4,500 6,000 3,600 15 16 Operating expenses for the year allocated to this project total $15,000. Lots unsold at the year-end were as follows Group 1 Group 2 Group 3 5 lots 7 lots 2 lots At the end of the fiscal year Crane Realty Corporation instructs you to arrive at the net income realized on this operation to date. (Round ratios for computational purposes to 4 decimal places, eg. 78.723490 and final answer to 0 decimal places, eg. 5,845.) Net income
Answer #1
Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.
Group No of Lots Price per lot Selling Price Relative Sales Price Total Cost Cost Allocated Cost per Lot Lot Sold Total Sale Total Cost of Goods Sold
Working 52000+30000
A B A*B=C D E E*D=F G=F/A H I=H*B J=H*G
                                              1                 9 $         4,500 $       40,500 40500/188100 21.53% $          82,000 $          17,656 $         1,962                4 $ 18,000 $                              7,847
                                              2               15 $         6,000 $       90,000 90000/188100 47.85% $          82,000 $          39,234 $         2,616                8 $ 48,000 $                            20,925
                                              3               16 $         3,600 $       57,600 57600/188100 30.62% $          82,000 $          25,110 $         1,569             14 $ 50,400 $                            21,971
$    188,100 $          82,000 $116,400 $                            50,743
Calculation of Net Income:
Sale $ 116,400
Less: Cost of Goods Sold $ -50,743
Gross Profit $   65,657
Less: Operating Expense $ -15,000
Net Income $   50,657